OpenStudy (howard-wolowitz):

Can someone help me here:

OpenStudy (phi):

You should not repost the same question. that is rude.

OpenStudy (anonymous):

I can Help!

OpenStudy (howard-wolowitz):

I just wonted to start fresh! Not intended to be rude!

OpenStudy (anonymous):

so 950$ 6.5% 10years right ok so make 6.5%=0.065+1=1.065 so 1.065 is your intreast in decimal form and you put 950 into thing and thing goes 1.065+950*10=9510.65 full anwsere

OpenStudy (howard-wolowitz):

no way

OpenStudy (howard-wolowitz):

@Haseeb96

OpenStudy (anonymous):

why no way?

OpenStudy (howard-wolowitz):

The answer cant go that high

OpenStudy (anonymous):

oh yes and round to the nearest dollar

OpenStudy (anonymous):

9511

OpenStudy (anonymous):

and yes it can be he put 950 into 10 years then intreast 6.5%

OpenStudy (howard-wolowitz):

so you think 9511 or 950 ?

OpenStudy (anonymous):

Face to the palm 950 was * by 10 years and + by the intreast 1.065 then rounded to the nearest dollar

OpenStudy (howard-wolowitz):

that seems right, I mean it is for 10 years 950

OpenStudy (anonymous):

ok well if you get it wrong its my fualt i really suck at math

OpenStudy (howard-wolowitz):

@iGreen can you help us here, by checking